a) Write the following numbers in scientific notation form. (i) 5333 (ii) 0.0025 (iii) 62 × 1010

(b) In 2003 Greg Murphy set the fastest lap ever on the Mount Panorama Circuit of length 6.213 km at Bathurst with a time of 126.86 seconds. (i) Calculate his average speed for this lap in m/s. (ii) What would be the average velocity? (iii) From a standing start, Greg reaches a velocity of 17.5 m/s in 5.0 s. Calculate the average acceleration during this time.
Answer #1

a) i) 5333 = 5.3 x 103

ii) 0.0025 = 2.5 x 10-3

iii) 62 x 1010 = 62620 = 6.2620 x 104

or  62 x 1010

b) average speed = (total distance / total time) = (6.213 x 103 m) / 126.86 s) = 48.975 m/s

average velocity = displacement / time = 0/126.86 = 0 m/s (displacement is zero where starting and ending points are equal)

average acceleration = change in velocity / change in time = 17.5 m/s / 5.0 s = 3.5 m/s2.